A man invests $10,000.00 in two enterprises. One enterprise yields 8% per year and the other yields 10% per year. If the income from both investments is $940.00 annually, how much did he invest in each enterprise?: A man invests $10,000.00 in two enterprises. One enterprise yields 8% per year and the other yields 10% per year. If the income from both investments is $940.00 annually, how much did he invest in each enterprise?
Answer by josmiceli(2018)

Let a= the 8% per yr investment
Let b= the 10% per year investment
.08a + .1b = 940
a + b = 10000
b = 10000 - a
.08a + .1*(10000 - a) = 940
.08a + 1000 - .1a = 940
-.02a = -60
a = 60/.02
a = 3000
And, since
a + b = 10000,
3000 + b = 10000
b = 7000
He invested $3,000 in the 8% investment
He invested $7,000 in the 10% investment
check:
.08a + .1b = 940
.08*3000 + .1*7000 = 940
240 + 700 = 940
940 = 940
OK